OpenStudy (anonymous):

does anyone know how to calculate expected value? i have some problems that i can't solve.

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

1000 chances are sold at $5 a piece for a raffle. Grand prize of $650, two prizes of $275, and 5 prizes of $100. first, calculate the expected value. if it is not a fair game, find the fair price for the game.

jimthompson5910 (jim_thompson5910):

E[ X ] = sum of probabilities * values E[ X ] = (1/1000)*650 + (2/1000)*275 + (5/1000)*100 + (992/1000)*(-5) E[ X ] = ???

OpenStudy (anonymous):

-3.26?

jimthompson5910 (jim_thompson5910):

you got it, so does that mean we have a fair game?

OpenStudy (anonymous):

no because fair game would be zero. so how exactly do I figure out the fair price?

jimthompson5910 (jim_thompson5910):

let p = price you want E[ X ] = 0 and you want to find the value of p that makes it zero

jimthompson5910 (jim_thompson5910):

E[ X ] = (1/1000)*650 + (2/1000)*275 + (5/1000)*100 + (992/1000)*p 0 = (1/1000)*650 + (2/1000)*275 + (5/1000)*100 + (992/1000)*p solve for p

jimthompson5910 (jim_thompson5910):

oh sry, it should be 0 = (1/1000)*650 + (2/1000)*275 + (5/1000)*100 + (992/1000)*(-p) solve for p to get your fair price

OpenStudy (anonymous):

-1.71?

OpenStudy (anonymous):

sorry i mean $1.71? per ticket

jimthompson5910 (jim_thompson5910):

good, I'm getting the same pretty much

jimthompson5910 (jim_thompson5910):

so if the price is $5 a ticket, then you expect to lose $3.26 on average if the price is around $1.71 a ticket, then you won't lose or gain any money. So it's a fair game for both parties.

OpenStudy (anonymous):

thanks! i get it now. i have two other questions, can you help me with those?

OpenStudy (anonymous):

a student is taking a standardized test consisting of multiple choice questions. one point is awarded for each correct answer. questions left blank do not get or lose points. if there are 8 options for each question and the student is penalized 1/2 point for each wrong answer, how many options must the student be able to rule out before the expected value of guessing is zero?

OpenStudy (anonymous):

is that question similar to the other?

jimthompson5910 (jim_thompson5910):

Does it say how many questions total there are?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

ok let me think for a sec

jimthompson5910 (jim_thompson5910):

I think it's E[ X ] = (1/8)*1 + (7/8)*(-1/2) but I'm still figuring out how to factor in the ones leave blank...hmm

jimthompson5910 (jim_thompson5910):

*you leave blank*

OpenStudy (anonymous):

i got -.3125 for that so far

jimthompson5910 (jim_thompson5910):

maybe it's a case where we don't leave any blank

jimthompson5910 (jim_thompson5910):

let's assume that

jimthompson5910 (jim_thompson5910):

so if n = number of options you rule out then 8 - n is the number of options left over 8 - n - 1 = 7 - n is the number of wrong options so this means E[ X ] = (1/(8-n))*1 + ((7-n)/(8-n))*(-1/2) 0 = (1/(8-n))*1 + ((7-n)/(8-n))*(-1/2) solve that for n to get your answer

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

im having trouble calculating it but it could be -5??

jimthompson5910 (jim_thompson5910):

it's 5 actually, so you lost a sign somewhere

jimthompson5910 (jim_thompson5910):

so if you rule out 5 options, then randomly guessing on every question you attempt will net you 0 points on average

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

i have one more

jimthompson5910 (jim_thompson5910):

this means that you will have 8-5 = 3 options to choose from and the expected value is E[ X ] = (1/3)*1 + (2/3)*(-1/2) E[ X ] = 0 which confirms the answer

OpenStudy (anonymous):

assume you have a laptop worth $2,500 and there is a 4% chance that it will be lost or stolen during the next year. what would be a fair premium for the insurance? (assuming the insurance provider does not make a profit)

OpenStudy (anonymous):

so the number of options the student must be able to rule out before the expected value of guessing becomes zero is 5?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok, thank you

jimthompson5910 (jim_thompson5910):

Let p = premium amount E[ X ] = (0.04)*(2500) + (0.96)*(-p) 0 = (0.04)*(2500) + (0.96)*(-p) ... Fair premium is when E[X] = 0 solve for p and round to the nearest cent

OpenStudy (anonymous):

so would i divide 100 by -0.96 to get p?

jimthompson5910 (jim_thompson5910):

correct, you would get 0 = (0.04)*(2500) + (0.96)*(-p) 0 = 100 + (0.96)*(-p) -100 = (0.96)*(-p) 0.96p = 100 p = 100/0.96 p = 104.166666666667 p = 104.17

jimthompson5910 (jim_thompson5910):

so the fair premium amount (for both parties) is $104.17

OpenStudy (anonymous):

okay, thanks so much for the help!

jimthompson5910 (jim_thompson5910):

np